# Cumulants in rectangular finite free probability and beta-deformed singular values

### v1

@@ 1.1 Preface
Finally, it is worth mentioning that there are other recent works [CDM-23, CD-24+] that study the quantized setting of $\beta$-tensor products of symmetric group representations and which have led to yet different and new one-parameter deformations of free convolution and cumulants. Whether those objects are also linked to Finite Free Probability is an open problem.
@@ 1.2 General setting
where $g(z):=\sum_{\ell=1}^{\infty}{\kappa^{\mathbf{a}}_{2\ell}z^{2\ell-1}}$; the operator $*_{g}$ is multiplication by $g(z)$; and $\partial_{t,u}$ is the operator defined on monomials by $\partial_{t,u}1:=0$, and $\partial_{t,u}z^{2m-1}:=(u+m-1)z^{2m-2}$, $\partial_{t,u}z^{2m}:=(t+m)z^{2m-1}$, for all $m\in\mathbb{Z}_{\geq 1}$.
@@ 1.3 Special setting: Rectangular convolution
This definition is motivated by random matrix theory. Specifically, if we consider independent $d\times(n+d)$ random matrices $A,B$ with deterministic singular values and $p(x)$, $r(x)$ are the characteristic polynomials of $AA^{*},BB^{*}$, then $(p\boxplus_{d}^{n}r)(x)$ is the expected characteristic polynomial of $(A+UBV)(A+UBV)^{*}$, where $U\in U(d)$ and $V\in U(n+d)$ are Haar-distributed; see [GM-22, MSS-22] for details. A related result is that if all roots of $p(x),r(x)$ are real and nonnegative, then the same is true of $(p\boxplus_{d}^{n}r)(x)$; see [GM-22, Thm. 2.3].
@@
Equivalently, this can be thought of as the empirical root distribution of the polynomial $\widetilde{p}(x):=p(x^{2})$ with roots ±$\pm\alpha_{1},\dots,\pm\alpha_{d}$.
@@
where $[z^{2\ell}]f(z)$ is the coefficient of $z^{2\ell}$ in the formal power series $f(z)\in\mathbb{R}[[z]]$. This definition is motivated by Eqn. (1), upon the specialization of parameters
@@
(b) Let $q\in[1,\infty)$, let $\{p_{d}(x)\}_{d\geq 1}$ be such that each $p_{d}(x)$ is a monic polynomial of degree $d$ with nonnegative real roots, and let $\mu$ be a symmetric [^3] A measure $\mu$ on $\mathbb{R}$ is said to be symmetric if $\mu(A)=\mu(-A)$, for all Lebesgue sets $A\subseteq\mathbb{R}$, where $-A:=\{-a\mid a\in A\}$. probability measure on the real line with finite moments of all orders. Then $\widetilde{\mu}[p_{d}]\to\mu$, as $d\to\infty$, in the sense of moments, if and only if
@@ 1.5 Organization of the paper
The reader might be curious as to why most sequences in this paper are labeled by the positive even integers (notably $(m_{2},m_{4},\dots)$ and $(\kappa_{2},\kappa_{4},\dots)$), as opposed to all positive integers. This boils down to the fact that [BG-09] and [Xu-23] chose to study the symmetric empirical root distributions (4) of polynomials with nonnegative real roots, as opposed to the empirical root distributions. Such choice implies that all measures that appear in their studies have vanishing odd moments $m_{1}=m_{3}=\cdots=0$ and vanishing odd cumulants $\kappa_{1}=\kappa_{3}=\cdots=0$. We decided to stick to the notations used by these authors, thus leading to the current presentation. If one considered instead the empirical root distributions, we could rewrite part of the theory for probability measures supported on $[0,\infty)$ instead; for them, generally, odd-indexed $m_{2n-1}$ and $\kappa_{2n-1}$ will be nonzero.
@@ Acknowledgments
The author would like to thank Jiaming Xu, Hoi Nguyen, Volodymyr Chub and Stanislav Surmylo for helpful conversations. This article is a follow-up of Chub-Surmylo [CS-24], which is the final report for a project from the research program Yulia’s Dream; the author is grateful to the organizers of that program for allowing him to play the role of research mentor in it. This work was partially supported by the NSF grant DMS-2348139.
@@ 2.1 Set partitions
Let $n\in\mathbb{Z}_{\geq 1}$. A set partition of $[n]:=\{1,2,\cdots,n\}$ is an unordered collection of pairwise disjoint nonempty subsets $B_{1},\cdots,B_{k}$ of $[n]$ such that $[n]=B_{1}\cup\cdots\cup B_{k}$; the corresponding set partition is denoted $\pi=\{B_{1},\cdots,B_{k}\}$. The subsets $B_{1},\cdots,B_{k}$ are called the blocks of $\pi$. The cardinalities of the blocks are denoted $|B_{1}|,\cdots,|B_{k}|$. Also, denote the number of blocks of $\pi$ by #$\#(\pi)$; in our running example, #$\#(\pi)=k$. Finally, the set of all set partitions of $[n]$ will be denoted by $\mathcal{P}(n)$.
@@
When $n=2n^{\prime}$ is even, we distinguish yet another class of set partitions. We say that $\pi\in\mathcal{P}(2n^{\prime})$ is even if all its blocks have even cardinalities. Denote the set of all even set partitions of $[2n^{\prime}]$ by $\mathcal{P}^{\textrm{even}}(2n^{\prime})\subseteq\mathcal{P}(2n^{\prime})$. Also, denote the the set of all even noncrossing set partitions of $[2n^{\prime}]$ by $\mathcal{NC}^{\textrm{even}}(2n^{\prime}):=\mathcal{P}^{\textrm{even}}(2n^{\prime})\cap\mathcal{NC}(2n^{\prime})$.
@@ 2.2 Generating series and set partitions
By definition, a Łukasiewicz path of length $n$ is a lattice path in $\mathbb{Z}^{2}$ from $(0,0)$ to $(n,0)$ that never goes below the $x$-axis and has $n$ steps, each being of the form $(1,j)$, for some $j\in\mathbb{Z}$, $j\geq-1$. A step $(1,j)$, for some $j\geq 1$ (resp. $j=0$ or $j=-1$) is called an up step (resp. horizontal or down step); also, if $(i,m)$ is the vertex of some Łukasiewicz path, we say that this vertex is at height $m$. We denote by $\mathbf{L}(n)$ the set of all Łukasiewicz paths of length $n$.
@@
Given $\pi=\{B_{1},\dots,B_{k}\}\in\mathcal{NC}(n)$, let $a_{i}\in[n]$ be the smallest element in $B_{i}$, for $1\leq i\leq k$. Next, construct a lattice path $\Lambda(\pi)$ that begins at $(0,0)$ and is followed by $n$ steps: its $m$-th step is $(1,|B_{i}|-1)$, if $m=a_{i}$ for some $i$, while it is $(1,-1)$ otherwise. Then the constructed lattice path $\Lambda(\pi)$ is a Łukasiewicz path of length $n$ and the map $\pi\mapsto\Lambda(\pi)$ is a bijection between $\mathcal{NC}(n)$ and $\mathbf{L}(n)$.
We say that a Łukasiewicz path is odd if all its steps are of the form $(1,2k-1)$, for some $k\in\mathbb{Z}_{\geq 0}$. Note that any odd Łukasiewicz path must have even length. If $n=2n^{\prime}$, then denote by $\mathbf{L}^{\textrm{odd}}(2n^{\prime})\subseteq\mathbf{L}(2n^{\prime})$ the set of odd Łukasiewicz paths of length $2n^{\prime}$. For example, some of the odd Łukasiewicz paths belonging to $\mathbf{L}^{\textrm{odd}}(6)$ are shown in Fig. 1 (ignore the step-labels in the figure, for now).
@@
$\bullet$ If the $m$-th step of $\Lambda(\pi)\in\mathbf{L}^{\textrm{odd}}(2n^{\prime})$ is a down step from some odd height $12s+$ 1 to the even height $2s$, then $m$ is even and is not the smallest element of any block of $\pi$.
$\bullet$ If the $m$-th step of $\Lambda(\pi)\in\mathbf{L}^{\textrm{odd}}(2n^{\prime})$ is a down step from some even height $2s$ to the odd height $12s-$ 1, then $m$ is odd and is not the smallest element of any block of $\pi$.
@@
[…]
@@ 3.2 In terms of operators applied to formal power series
the operator $*_{g}$ is multiplication by $g(z)$, and $[z^{0}]\colon\mathbb{R}[[z]]\to\mathbb{R}$ picks up the constant term of a power series, i.e. $[z^{0}]f(z)=f(0)$.
@@
We first find out how $d_{t,u}$ acts on certain products:
$d_{t,u}(z^{2a+1}z^{2b})$ $=d_{t,u}(z^{2a+2b+1})=\left(u-\frac{1}{2}\right)\cdot z^{2a+2b}=(d_{t,u}z^{2a+1})\cdot z^{2b},\,\text{ if $a,b\in\mathbb{Z}_{\geq 0}$},$
$d_{t,u}(z^{2a}z^{2b})$ $=d_{t,u}(z^{2a+2b})=t\cdot z^{2a+2b-1}=(d_{t,u}z^{2a})\cdot z^{2b},\,\text{ if $a\in\mathbb{Z}_{\geq 1},b\in\mathbb{Z}_{\geq 0}$}.$
Note that the last equality is false if $a=0$, but $d_{t,u}(1\cdot z^{2b})=\mathbf{1}_{\{b\neq 0\}}\cdot tz^{2b-1}=1\cdot d_{t,u}z^{2b}$, for all $b\in\mathbb{Z}_{\geq 0}$. By linearity of the operator $d_{t,u}$, the previous equations imply
@@
Step 4 (Conclusion of the argument). As mentioned in the last sentence of Step 2, to finish the proof of the theorem, it will suffice to show that the $a_{2n}$’s defined by (14) and $m_{2k}$’s defined by (16) satisfy the recursive relations (15). Note that $g(z)=\sum_{\ell\geq 1}{\kappa_{2\ell}z^{2\ell-1}}$ only has odd powers of $z$, so $[z^{0}](g(z))=0$. Then it follows from Claim 3.1 that
@@
We want to take the constant term to both sides of this equation. For the right hand side, we use that $[z^{0}](u(z)v(z))=[z^{0}](u(z))\cdot[z^{0}](v(z))$, for any $u,v\in\mathbb{R}[[z]]$, and also that $[z^{0}]e^{G(z)}=1$, because $G(z)$ does not have constant terms. Then (22) yields
$[z^{0}]\,\partial_{t,u}\big{(}g(z)e^{G(z)}\big{)}=[z^{0}](\partial_{t,u}+*_{g})(g(z)).$
@@
Now note that $\partial_{t,u}$ decreases the degree of a power of $z$ by $1$. Also, because $g(z)$ has only odd powers of $z$, it follows that $(\partial_{t,u}+*_{g})(g(z))$ only has even powers of $z$, and $(\partial_{t,u}+*_{g})^{2}(g(z))$ only has odd powers of $z$; in particular, $[z^{0}]\,(\partial_{t,u}+*_{g})^{2}(g(z))=0$. Now we can apply $\partial_{t,u}$ to (23) and use Claim 3.1 again to get
@@
$[z^{0}]\,\partial_{t,u}^{2n-1}(g(z)e^{G(z)})=[z^{0}]\,(\partial_{t,u}+*_{g})^{2n-1}(g(z))\\ +\sum_{m=1}^{n-1}{[z^{0}](\partial_{t,u}+*_{g})^{2m-1}(g(z))\cdot[z^{0}]\,\partial_{t,u}^{2n-2m-1}d_{t,u}(e^{G(z)})}.$ (26)
To simplify $[z^{0}]\,\partial_{t,u}^{2n-2m-1}d_{t,u}(e^{G(z)})$, we use the following obvious trick.
Trick 3.1.
If $\partial_{1},\dots,\partial_{k}$ are operators on $\mathbb{R}[[z]]$, each of which is a multiple of $\frac{d}{dz}$, then
$[z^{0}]\,\partial_{1}\cdots\partial_{k}(f(z))=\partial_{1}\cdots\partial_{k}\big{(}[z^{k}]\,f(z)\big{)}.$
Since $\partial_{t,u}$ and $d_{t,u}$ are multiples of $\frac{d}{dz}$, we deduce
$[z^{0}]\,\partial_{t,u}^{2n-2m-1}d_{t,u}(e^{G(z)})$ $=\partial_{t,u}^{2n-2m-1}d_{t,u}\big{(}[z^{2n-2m}]e^{G(z)}\big{)}$
$=\partial_{t,u}^{2n-2m-1}\left(\frac{t}{2n-2m}\cdot\frac{d}{dz}\right)\big{(}[z^{2n-2m}]e^{G(z)}\big{)}$
@@
where the first and third equalities follow from the Trick 3.1, the second equality holds because $d_{t,u}z^{2\ell}=\frac{t}{2\ell}\cdot\frac{d}{dz}(z^{2\ell})$, by the definition of the action of $d_{t,u}$ and the last equality is because $\frac{G^{\prime}(z)}{2}=g(z)$. Plugging (27) into (26) gives:
$[z^{0}]\,\partial_{t,u}^{2n-1}(g(z)e^{G(z)})=[z^{0}]\,(\partial_{t,u}+*_{g})^{2n-1}(g(z))\\ +\sum_{m=1}^{n-1}{[z^{0}](\partial_{t,u}+*_{g})^{2m-1}(g(z))\cdot\left(\frac{t}{n-m}\cdot[z^{0}]\,\partial_{t,u}^{2n-2m-1}(g(z)e^{G(z)})\right)},$ (28)

### v3

@@ 1.1 Preface
Finally, it is worth mentioning that there are other recent works [CDM-23, CD-25a, CD-25b] that study the quantized setting of $\beta$-tensor products of symmetric group representations and which have led to yet different and new one-parameter deformations of free convolution and cumulants. Whether those objects are also linked to Finite Free Probability is an open problem.
@@ 1.2 General setting
where $g(z):=\sum_{\ell=1}^{\infty}{\kappa^{\mathbf{a}}_{2\ell}z^{2\ell-1}}$; the operator $_{g}$ is multiplication by $g(z)$; and $\partial_{t,u}$ is the operator defined on monomials by $\partial_{t,u}1:=0$, and $\partial_{t,u}z^{2m-1}:=(u+m-1)z^{2m-2}$, $\partial_{t,u}z^{2m}:=(t+m)z^{2m-1}$, for all $m\in\mathbb{Z}_{\geq 1}$.
@@ 1.3 Special setting: Rectangular convolution
Alternatively, the $(n,d)$-rectangular convolution can be defined as follows. If $D:=\frac{d}{dx}$, there are unique polynomials $P(x)$, $R(x)$ of degree $d$ such that $p(x)=P(x^{-n}Dx^{n+1}D)x^{d}$ and $r(x)=R(x^{-n}Dx^{n+1}D)x^{d}$; then $(p\boxplus^{n}_{d}r)(x)=P(x^{-n}Dx^{n+1}D)R(x^{-n}Dx^{n+1}D)x^{d}$. The case $n=0$ is an observation of Mirabelli [Mir-21], also stated in [MSS-22]. This crucial fact will be used in the proof of our Main Result III on the asymptotic theory of nonnegative real-rooted polynomials and, in fact, it was essential in the conception of its statement.
This convolution is motivated by random matrix theory. Specifically, if we consider independent $d\times(n+d)$ random matrices $A,B$ with deterministic singular values and $p(x)$, $r(x)$ are the characteristic polynomials of $AA^{*},BB^{*}$, then $(p\boxplus_{d}^{n}r)(x)$ is the expected characteristic polynomial of $(A+UBV)(A+UBV)^{*}$, where $U\in U(d)$ and $V\in U(n+d)$ are Haar-distributed; see [GM-22, MSS-22] for details.
A last related result that we will mention is that if all roots of $p(x),r(x)$ are real and nonnegative, then the same is true of $(p\boxplus_{d}^{n}r)(x)$; see [GM-22, Thm. 2.3].
@@
Equivalently, this can be thought of as the empirical root distribution of the polynomial $\widetilde{p}(x):=p(x^{2})$ with roots $\pm\alpha_{1},\dots,\pm\alpha_{d}$.
@@
where $z^{2\ell}]f(z)$ is the coefficient of $z^{2\ell}$ in the formal power series $f(z)\in\mathbb{R}[[z]]$. This definition is motivated by Eqn. (1), upon the specialization of parameters
@@
(b) Let $q\in[1,\infty)$, let $\{p_{d}(x)\}_{d\geq 1}$ be such that each $p_{d}(x)$ is a monic polynomial of degree $d$ with nonnegative real roots, and let $\mu$ be a symmetric [^3] A measure $\mu$ on $\mathbb{R}$ is said to be symmetric if $\mu(A)=\mu(-A)$, for all Lebesgue sets $A\subseteq\mathbb{R}$, where $A:=\{-a\mid a\in A\}$. probability measure on the real line with finite moments of all orders. Then $\widetilde{\mu}[p_{d}]\to\mu$, as $d\to\infty$, in the sense of moments, if and only if
@@ 1.5 Organization of the paper
The reader might be curious as to why most sequences in this paper are labeled by the positive even integers (notably $(m_{2},m_{4},\dots)$ and $(\kappa_{2},\kappa_{4},\dots)$), as opposed to all positive integers. This boils down to the fact that [BG-09] and [Xu-23] chose to study the symmetric empirical root distributions (4) of polynomials with nonnegative real roots, as opposed to the empirical root distributions. Such choice implies that all measures that appear in their studies have vanishing odd moments $m_{1}=m_{3}=\cdots=0$ and vanishing odd cumulants $\kappa_{1}=\kappa_{3}=\cdots=0$. We decided to stick to the notations used by these authors, thus leading to the current presentation. If one considered instead the empirical root distributions, we could rewrite part of the theory for probability measures supported on $0,\infty)$ instead; for them, generally, odd-indexed $m_{2n-1}$ and $\kappa_{2n-1}$ will be nonzero.
@@ Acknowledgments
The author would like to thank Jiaming Xu, Hoi Nguyen, Volodymyr Chub and Stanislav Surmylo for helpful conversations, as well as two anonymous referees for helpful comments. This article is a follow-up to the final report by Chub–Surmylo of a project from the research program Yulia’s Dream; the author is grateful to the organizers for allowing him to play the role of research mentor. This work was partially supported by the NSF grant DMS-2348139.
@@ 2.1 Set partitions
Let $n\in\mathbb{Z}_{\geq 1}$. A set partition of $n]:=\{1,2,\cdots,n\}$ is an unordered collection of pairwise disjoint nonempty subsets $B_{1},\cdots,B_{k}$ of $[n]$ such that $n]=B_{1}\cup\cdots\cup B_{k}$; the corresponding set partition is denoted $\pi=\{B_{1},\cdots,B_{k}\}$. The subsets $B_{1},\cdots,B_{k}$ are called the blocks of $\pi$. The cardinalities of the blocks are denoted $|B_{1}|,\cdots,|B_{k}|$. Also, denote the number of blocks of $\pi$ by $(\pi)$; in our running example, $(\pi)=k$. Finally, the set of all set partitions of $[n]$ will be denoted by $\mathcal{P}(n)$.
@@
When $n=2n^{\prime}$ is even, we distinguish yet another class of set partitions. We say that $\pi\in\mathcal{P}(2n^{\prime})$ is even if all its blocks have even cardinalities. Denote the set of all even set partitions of $2n^{\prime}]$ by $\mathcal{P}^{\textrm{even}}(2n^{\prime})\subseteq\mathcal{P}(2n^{\prime})$. Also, denote the the set of all even noncrossing set partitions of $2n^{\prime}]$ by $\mathcal{NC}^{\textrm{even}}(2n^{\prime}):=\mathcal{P}^{\textrm{even}}(2n^{\prime})\cap\mathcal{NC}(2n^{\prime})$.
@@ 2.2 Generating series and set partitions
By definition, a Łukasiewicz path of length $n$ is a lattice path in $\mathbb{Z}^{2}$ from $(0,0)$ to $(n,0)$ that never goes below the $x$-axis and has $n$ steps, each being of the form $(1,j)$, for some $j\in\mathbb{Z}$, $j\geq-1$. A step $(1,j)$, for some $j\geq 1$ (resp. $j=0$ or j=1j=-1) is called an up step (resp. horizontal or down step); also, if $(i,m)$ is the vertex of some Łukasiewicz path, we say that this vertex is at height $m$. We denote by $\mathbf{L}(n)$ the set of all Łukasiewicz paths of length $n$.
@@
Given $\pi=\{B_{1},\dots,B_{k}\}\in\mathcal{NC}(n)$, let $a_{i}\in[n]$ be the smallest element in $B_{i}$, for $1\leq i\leq k$. Next, construct a lattice path $\Lambda(\pi)$ that begins at $(0,0)$ and is followed by $n$ steps: its $m$-th step is $(1,|B_{i}|-1)$, if $m=a_{i}$ for some $i$, while it is (1,1)(1,-1) otherwise. Then the constructed lattice path $\Lambda(\pi)$ is a Łukasiewicz path of length $n$ and the map $\pi\mapsto\Lambda(\pi)$ is a bijection between $\mathcal{NC}(n)$ and $\mathbf{L}(n)$.
We say that a Łukasiewicz path is odd if all its steps are of the form (1,2k1)(1,2k-1), for some $k\in\mathbb{Z}_{\geq 0}$. Note that any odd Łukasiewicz path must have even length. If $n=2n^{\prime}$, then denote by $\mathbf{L}^{\textrm{odd}}(2n^{\prime})\subseteq\mathbf{L}(2n^{\prime})$ the set of odd Łukasiewicz paths of length $2n^{\prime}$. For example, some of the odd Łukasiewicz paths belonging to $\mathbf{L}^{\textrm{odd}}(6)$ are shown in Fig. 1 (ignore the step-labels in the figure, for now).
@@
$\bullet$ If the $m$-th step of $\Lambda(\pi)\in\mathbf{L}^{\textrm{odd}}(2n^{\prime})$ is a down step from some odd height $2s+1$ to the even height $2s$, then $m$ is even and is not the smallest element of any block of $\pi$.
$\bullet$ If the $m$-th step of $\Lambda(\pi)\in\mathbf{L}^{\textrm{odd}}(2n^{\prime})$ is a down step from some even height $2s$ to the odd height 2s12s-1, then $m$ is odd and is not the smallest element of any block of $\pi$.
@@
There exists a symmetric probability measure, uniquely determined by its moments and with only its second $q$-rectangular free cumulant being nonzero. This is the $q$-rectangular analogue of the centered Gaussian distribution for the theory of $q$-rectangular free probability. It will be denoted by $\lambda^{(q)}_{\sigma^{2}}$, if it has variance $\sigma^{2}>0$. The density of $\lambda^{(q)}_{\sigma^{2}}$, stated next, was obtained in [BG-07, Thm. 4.3] for $\sigma^{2}=1$; see also [BG-09, Sec. 3.10.2] for general $\sigma^{2}>0$. [^5] There is a small typo in the formula from [BG-09, Sec. 3.10.2].
@@ 3.2 In terms of operators applied to formal power series
the operator $_{g}$ is multiplication by $g(z)$, and $z^{0}]\colon\mathbb{R}[[z]]\to\mathbb{R}$ picks up the constant term of a power series, i.e. $z^{0}]f(z)=f(0)$.
@@
Observe that $d_{t,u}$ acts as follows on certain products of powers of $z$:
$d_{t,u}(z^{2a+1}z^{2b})$ $=d_{t,u}(z^{2a+2b+1})=\left(u-\frac{1}{2}\right)\cdot z^{2a+2b}=(d_{t,u}z^{2a+1})\cdot z^{2b},\,\text{ for all $a,b\in\mathbb{Z}_{\geq 0}$},$
$d_{t,u}(z^{2a}z^{2b})$ $=d_{t,u}(z^{2a+2b})=t\cdot z^{2a+2b-1}=(d_{t,u}z^{2a})\cdot z^{2b},\,\text{ for all $a\in\mathbb{Z}_{\geq 1},b\in\mathbb{Z}_{\geq 0}$},$
$d_{t,u}(1\cdot z^{2b})$ $=\mathbf{1}_{\{b\neq 0\}}\cdot tz^{2b-1}=1\cdot d_{t,u}z^{2b},\,\text{ for all $b\in\mathbb{Z}_{\geq 0}$}.$
By linearity of the operator $d_{t,u}$, the previous equations imply
@@
Step 4 (Conclusion of the argument). As mentioned in the last sentence of Step 2, to finish the proof of the theorem, it will suffice to show that the $a_{2n}$’s defined by (14) and $m_{2k}$’s defined by (16) satisfy the recursive relations (15). Note that $g(z)=\sum_{\ell\geq 1}{\kappa_{2\ell}z^{2\ell-1}}$ only has odd powers of $z$, so $z^{0}](g(z))=0$. Then it follows from Claim 3.1 that
@@
We want to take the constant term to both sides of this equation. For the right hand side, we use that $z^{0}](u(z)v(z))=[z^{0}](u(z))\cdot[z^{0}](v(z))$, for any $u,v\in\mathbb{R}[[z]]$, and also that $z^{0}]e^{G(z)}=1$, because $G(z)$ does not have constant terms. Then (22) yields
$z^{0}]\,\partial_{t,u}\big(g(z)e^{G(z)}\big)=[z^{0}](\partial_{t,u}+*_{g})(g(z)).$
@@
Now note that $\partial_{t,u}$ decreases the degree of a power of $z$ by $1$. Also, because $g(z)$ has only odd powers of $z$, it follows that $(\partial_{t,u}+*_{g})(g(z))$ only has even powers of $z$, and $(\partial_{t,u}+*_{g})^{2}(g(z))$ only has odd powers of $z$; in particular, $z^{0}]\,(\partial_{t,u}+*_{g})^{2}(g(z))=0$. Now we can apply $\partial_{t,u}$ to (23) and use Claim 3.1 again to get
@@
$z^{0}]\,\partial_{t,u}^{2n-1}(g(z)e^{G(z)})=[z^{0}]\,(\partial_{t,u}+*_{g})^{2n-1}(g(z))\\ +\sum_{m=1}^{n-1}{[z^{0}](\partial_{t,u}+*_{g})^{2m-1}(g(z))\cdot[z^{0}]\,\partial_{t,u}^{2n-2m-1}d_{t,u}(e^{G(z)})}.$ (26)
To simplify $z^{0}]\,\partial_{t,u}^{2n-2m-1}d_{t,u}(e^{G(z)})$, we use the following obvious fact.
Fact 3.1.
Let $k\in\mathbb{Z}_{\geq 1}$ and let $\partial_{1},\dots,\partial_{k}$ be operators on $\mathbb{R}[[z]]$, each of which is a multiple of $\frac{d}{dz}$. For any $f(z)=\sum_{n=0}^{\infty}{a_{n}z^{n}}\in\mathbb{R}[[z]]$, if we denote $z^{k}]f(z):=a_{k}$, then we have
$z^{0}]\,\partial_{1}\cdots\partial_{k}(f(z))=\partial_{1}\cdots\partial_{k}\Big([z^{k}]f(z)\cdot z^{k}\Big).$
Since $\partial_{t,u}$ and $d_{t,u}$ are multiples of $\frac{d}{dz}$, we can then deduce
$[z^{0}]\,\partial_{t,u}^{2n-2m-1}d_{t,u}(e^{G(z)})$ $=\partial_{t,u}^{2n-2m-1}d_{t,u}\Big([z^{2n-2m}]e^{G(z)}\cdot z^{2n-2m}\Big)$
$=\partial_{t,u}^{2n-2m-1}\left(\frac{t}{2n-2m}\cdot\frac{d}{dz}\right)\Big([z^{2n-2m}]e^{G(z)}\cdot z^{2n-2m}\Big)$
@@
where the first and third equalities follow from Fact 3.1, the second equality holds because $d_{t,u}z^{2\ell}=\frac{t}{2\ell}\cdot\frac{d}{dz}(z^{2\ell})$, by the definition of the action of $d_{t,u}$ and the last equality is because $\frac{G^{\prime}(z)}{2}=g(z)$. Plugging (27) into (26) gives:
$z^{0}]\,\partial_{t,u}^{2n-1}(g(z)e^{G(z)})=[z^{0}]\,(\partial_{t,u}+*_{g})^{2n-1}(g(z))\\ +\sum_{m=1}^{n-1}{[z^{0}](\partial_{t,u}+*_{g})^{2m-1}(g(z))\cdot\left(\frac{t}{n-m}\cdot[z^{0}]\,\partial_{t,u}^{2n-2m-1}(g(z)e^{G(z)})\right)},$ (28)